# H3AE: High Compression, High Speed, and High Quality AutoEncoder for Video Diffusion Models

### v1

@@ 5 Conclusion
In this work, we systematically examine autoencoder architecture design and optimize the computation distribution to obtain a series of efficient AEs that can decode latents in real-time on mobile-device. With the ultra high spatial-temporal compression ratio, we successfully reduce the latent tokens and achieve faster generation speed for the DiT-based video diffusion model. We unify plain reconstruction and image-conditioned I2V reconstruction, demonstrate improved results for both settings with a single set of weight, simplifying applications and saving the potential adaptation cost. Our empirical study also shows that popular discriminative losses, i.e., GAN, LPIPS, and DWT losses, provide no significant improvement when training AEs at scale. We propose a novel latent consistency loss that requires no complicated discriminator design or hyperparameter tuning but provides stable improvements in reconstruction quality.
Appendix
@@
Table 7: Despite H3AE is designed as video autoencoder, it can also be used as an image autoencoder. The quality comparison of image reconstruction between current SOTA high-compression autoencoder i.e. DC-AE[9] and our H3AE on DAVIS and OpenVid-HD datasets. The results demonstrate H3AE outperforms DC-AE under same compression ratios.
Figure 6: More examples of videos generated by a 2B DiT denoiser, trained on the latent space of our H3AE.

### v2

@@ 5 Conclusion
In this work, we systematically examine autoencoder architecture design and optimize the computation distribution to obtain a series of efficient AEs that can decode latents in real-time on mobile device. With the ultra high spatial-temporal compression ratio, we successfully reduce the latent tokens and achieve faster generation speed for the DiT-based video diffusion model. We unify plain reconstruction and image-conditioned I2V reconstruction, demonstrating improved results for both settings with a single set of weights, simplifying applications and saving the potential adaptation cost. Our empirical study also shows that popular discriminative losses, i.e., GAN, LPIPS, and DWT losses, provide no significant improvement when training AEs at scale. We propose a novel latent consistency loss that does not require complicated discriminator design or hyperparameter tuning but provides stable improvements in reconstruction quality. We discuss limitations and broader impact in Appendix B.
Appendix A Use of LLMs
Large language models (e.g., ChatGPT, Gemini) were used exclusively for grammar polishing and formatting assistance. All proposed concepts, experiment design, and analysis are NOT generated by LLMs.
Appendix B Limitations and Broader Impact
While H3AE advances the design and training of video VAEs, several limitations remain. Though we demonstrate mobile deployment, our evaluations are still bounded by current hardware and model scales; larger-scale models may reveal new bottlenecks. Additionally, both VAE and DiT training assume access to large-scale video datasets, which are not universally available. We hope that this can be standardized for the research community in the near future.
For the broader impact of H3AE, efficient VAE makes high-quality video generation more accessible by reducing the computational and memory demands of diffusion models. This democratizes research and creative applications, enabling use on consumer devices and in resource-constrained environments. However, the same accessibility raises concerns about misuse, such as the large-scale generation of misleading or harmful content.
Appendix C H3AE as Image Autoencoder
The H3AE can be used as an image VAE and its image reconstruction performance is presented in Tab. 7. Although H3AE targets a video autoencoder, it can still be utilized as an image autoencoder. We compare H3AE’s performance with current SOTA high-compression autoencoder DC-AE (Chen et al., 2025b), showing that H3AE outperforms DC-AE under the same compression ratios Tab. 7.
Table 7: Despite H3AE is designed as video autoencoder, it can also be used as an image autoencoder. The quality comparison of image reconstruction between current SOTA high-compression autoencoder i.eChen et al. (2025b) and our H3AE on DAVIS and OpenVid-HD datasets. The results demonstrate H3AE outperforms DC-AE under same compression ratios.
@@ Appendix C H3AE as Image Autoencoder
Appendix D H3AE as Image-to-Video Autoencoder
Figure 5 demonstrates the performance of our omni-objective training in the image-to-video (I2V) setting. By randomly dropping the first-frame condition during training, H3AE learns to flexibly decode with or without image guidance. As shown, conditioning on the first frame enables the VAE to better preserve details, while still maintaining robustness when the condition is absent. This demonstrates that omni training not only unifies T2V and I2V within a single model but also enhances reconstruction quality in both regimes.
Appendix E Experiment fairness.
To eliminate the impact of different data distributions and ensure fair comparisons, we re-trained the LTX VAE (HaCohen et al., 2024) on the same dataset and under the same training setup as H3AE, as in Tab. 8. We find that with our video dataset and training recipe, similar results are obtained compared to the official LTX VAE weights. Due to different training stages and loss weights, our retrained LTX VAE is slightly better at reconstruction metrics but a bit worse in rFVD. Due to the reconstruction nature, Video VAE training does not rely a lot on video data distribution and quality. Our dataset and training strength is on par with public Video VAE works and thus produce comparable results.
Table 8: Dataset ablation.
[…]
LTX-VAE $8\times 32\times 32$ 28.34 0.7923 165.43
LTX-VAE-Ours $8\times 32\times 32$ 28.78 0.8133 185.33
H3AE $8\times 32\times 32$ 30.23 0.8412 122.82
Appendix F Model Architecture.
We provide the details of the optimized H3AE architecture in Tab. 9. Note that the spatial stage only handles spatial downsample and upsample, and the spatial-temporal stage applies either or both spatial and temporal sampling according to the designated configuration. For instance, for the $8\times 32\times 32$ setting, there are two $1\times 2\times 2$ downsamplings in the spatial stage and three $2\times 2\times 2$ downsamplings in the spatial-temporal stage.
Table 9: H3AE architecture reported with channel dimension and number of blocks. Note that there are two convolution layers in a residual block. S and ST refer to spatial and spatial-temporal stages, respectively. We use 8 heads for the causal attention.
Compression Ratio S Stage-1 Conv2D S Stage-2 Conv2D ST Stage-1 Conv3D ST Stage-2 Conv3D ST Stage-3 Conv3D ST Stage Attn.
$4\times 16\times 16$ $32\times 2$ $64\times 2$ $128\times 2$ $256\times 8$ - -
$8\times 32\times 32$ $32\times 2$ $64\times 2$ $128\times 2$ $256\times 6$ - $512\times 8$
$8\times 64\times 64$ $32\times 2$ $64\times 2$ $128\times 2$ $256\times 6$ $512\times 8$ $512\times 8$
Appendix G More Video Visualizations
We provide more video visualizations of the diffusion transformer trained under H3AE latent space in Fig. 6 and the supplementary files.
Figure 6: More videos generated by a 2B DiT denoiser, trained on the latent space of our H3AE.